Westerville Company reported the following results from last year’s operations:

Sales $ 1,800,000
Variable expenses 435,000
Contribution margin 1,365,000
Fixed expenses 1,005,000
Net operating income $ 360,000
Average operating assets $ 1,200,000

At the beginning of this year, the company has a $300,000 investment opportunity with the following cost and revenue characteristics:

Sales $ 360,000
Contribution margin ratio 70 % of sales
Fixed expenses $ 216,000

The company’s minimum required rate of return is 10%.

6. What is the ROI related to this year’s investment opportunity? (Do not round intermediate calculations.)

7. If the company pursues the investment opportunity and otherwise performs the same as last year, what margin will it earn this year? (Round your percentage answer to 1 decimal place (i.e., 0.1234 should be entered as 12.3).)

8. If the company pursues the investment opportunity and otherwise performs the same as last year, what turnover will it earn this year? (Round your answer to 2 decimal places.)

9. If the company pursues the investment opportunity and otherwise performs the same as last year, what ROI will it earn this year? (Do not round intermediate calculations. Round your percentage answer to 1 decimal place (i.e., 0.1234 should be considered as 12.3).)

10-a. If Westerville’s chief executive officer will earn a bonus only if her ROI from this year exceeds her ROI from last year, would she pursue the investment opportunity?

  • Yes

  • No

10-b. Would the owners of the company want her to pursue the investment opportunity?

  • Yes

  • No
Answer #1

6.ROI

net operating income/investment

Margin = 360000*70% - 216000

=252000-216000

=36000$

ROI = 36000/300000

=12%

7.

company new opportunity Total
sales 1,800,000 360,000 2,160,000
var exps 435000 108000[360000*30%] 543000
contribution margin 1365000 252000 1617000
fixed 1005000 216000 1221000
net operating income 360000 36000 396000

NOI / SALES

MARGIN = 396000/2160000

=18.3%

(8)TURNOVER

=sales/average operating asset

=2,160,000/ (1,200,000+300,000)

=2,160,000/1,500,000

=1.5

(9)ROI

=MARGIN *TURNOVER

=18.333333*1.44

=26.4%

(10)A

NO

as the ROI dropped from 30% to 12%

10b

yes

As the ROI is 12% which is higher than required rate of return that is 10%